# Multiscale Growth Kinetics of Model Biomolecular Condensates Under Passive and Active Conditions

### v1

@@ Abstract
Living cells exhibit a complex organization comprising numerous compartments, among which are RNA- and protein-rich membraneless, liquid-like organelles known as biomolecular condensates. Energy-consuming processes regulate their formation and dissolution, with (de-)phosphorylation by specific enzymes being among the most commonly involved reactions. By employing a model system consisting of a phosphorylatable peptide and homopolymeric RNA, we elucidate how enzymatic activity modulates the growth kinetics and alters the local structure of biomolecular condensates. Under passive condition, time-resolved ultra-small-angle X-ray scattering with synchrotron source reveals a nucleation-driven coalescence mechanism maintained over four decades in time, similar to the coarsening of simple binary fluid mixtures. Coarse-grained molecular dynamics simulations show that peptide-decorated RNA chains assembled shortly after mixing constitute the relevant subunits. In contrast, actively-formed condensates initially display a local mass fractal structure, which gradually matures upon enzymatic activity before condensates undergo coalescence. Both types of condensate eventually reach a steady state but fluorescence recovery after photobleaching indicates a peptide diffusivity twice higher in actively-formed condensates consistently with their loosely-packed local structure. We expect multiscale, integrative approaches implemented with model systems to link effectively the functional properties of membraneless organelles to their formation and dissolution kinetics as regulated by cellular active processes.
@@ 1 Introduction
Membraneless organelles are intracellular compartments devoid of lipid membrane, consisting of multiple proteins and possibly nucleic acids (DNA and/or RNA), and formed by liquid-liquid phase separation 1, 2, 3, 4. These biomolecular condensates perform essential biological functions by sequestering specific factors and promoting biochemical reactions following the needs of the cell 5. They also serve as protocells for the study of the emergence of functional cellular life or as catalytic microcompartments 6, 7. Many condensates found in eucaryotic cells contain RNA and are involved in gene expression, such as nucleoli 8, P granules 9, stress granules 10 or Cajal bodies 11 to name a few. Protein-RNA condensates may also play a role in pathological processes 12, for example by facilitating the transcription of viral RNA during SARS-Cov-2 replication 13. Membraneless organelles are relatively short-lived, in that they are formed to perform specific functions and are dissolved after these tasks are completed. Such a flexibility is maintained through energy-consuming enzymatic processes. In particular, various post-translational modifications have been shown to regulate intracellular phase separation, including phosphorylation 14, methylation 15, acetylation and combinations thereof 16.
@@
Phase separation is a generic phenomenon by which a homogeneous material, spontaneously or not, separates into two or more distinct domains (phases) with different physical and/or chemical properties. Extensive experimental and theoretical efforts have been devoted to elucidating the phase-separation kinetics of binary systems 17, including fluid mixtures relevant to biomolecular condensates, under both passive (spontaneous) and active (chemically driven) conditions 18, 19, 20. A major limitation in these investigations lies in the lack of information across length and time scales spanning several orders of magnitude. Indeed, biomolecular condensates are mostly micrometer-sized and therefore, fluorescence microscopy has become the standard technique to monitor their life cycle, especially within cells 8. Quite recently, holographic microscopy has demonstrated its ability to monitor the growth of PopZ-cation condensates without the need for protein labeling 21. In all cases, submicrometer-sized condensates formed along the growth pathway are overlooked, and events occurring on subsecond time scales often remain inaccessible. Conversely, coarse-grained molecular dynamics simulations can provide a wealth of information at the molecular level 22, and a few phase-separated protein-RNA condensates have been elucidated 23, 24. Nonetheless, coarse-grained models face substantial challenges in quantitatively reproducing the thermodynamic properties of biomolecular condensates. In addition, although they allow the simulation of relatively large supramolecular systems, the accessible time scales remain limited to a few microseconds at most.
@@
To bridge these multiple scales, time-resolved small-angle X-ray scattering (TR-SAXS) using synchrotron source has emerged as a state-of-the-art technique, enabling to track self-assembly processes with millisecond temporal resolution and spatial sensitivity ranging from a few to several hundred nanometers 25, 26, 27, 28, 29. Within the context of liquid-liquid phase separation, the assembly kinetics of a prion-like domain has revealed the existence of a double nucleation barrier to overcome before the condensate growth can proceed 30. With the advent of ultra-small angles available at a very few synchrotron facilities worldwide, time-resolved ultra-small-angle X-ray scattering (TR-USAXS) 31 has extended the capabilities of X-ray scattering to the micrometer scale, and the spinodal decomposition of bovine serum albumin induced by YCl3 was probed during the first two minutes following the mixing of components 32. Likewise, coacervates composed of oppositely-charged polyelectrolytes were found to change from mass fractal to surface fractal morphologies while growing by coalescence 33. Nevertheless, how enzymatic activity regulates the growth mechanism, the local structure, and the morphology of biomolecular condensates – compared to conditions lacking active processes – remains experimentally unaddressed.
@@
To fill this knowledge gap, we employ a model system of biomolecular condensates whose formation is triggered by the dephosphorylation of serine residues, thereby mimicking the enzymatic regulation encountered in intracellular membraneless organelles 34. This system consists of an arginine-rich peptide (RRASLRRASL) containing phosphorylatable serines, associated with poly(uridylic acid) (polyU) RNA. Condensates form spontaneously when the serines bear positive charges – a condition hereafter referred to as passive – or are enzymatically induced by lambda phosphatase (LPP), which removes phosphate groups from initially phosphorylated serines, defining the active condition. This flexible system enables the comparison between condensate formation mechanisms under passive and active conditions. We therefore combine TR-USAXS experiments performed at varying peptide and enzyme concentrations with confocal microscopy and coarse-grained molecular dynamics simulations to capture these mechanisms across a wide range of length and time scales.
@@ 2.1 Phase separation occurs in the cationic regime
Upon mixing RRASLRRASL peptide with anionic polyU RNA (Fig. 1a) in 4 mM MgCl2, 2.25 mM ATP, 1.6 mM dithiothreitol, 50 mM HEPES pH 7.4 at 37 °C, biomolecular condensates in the form of liquid-like droplets are spontaneously made by coacervation 34. A direct evidence for the presence of micrometer-sized droplets is given by the turbid aspect of the solution (Fig. 1b). Confocal microscopy images performed with 0.5% of fluorescent TAMRA-peptide confirm the spherical shape and the micrometer-scaled diameter of the condensates (Fig. 1c). At a given RNA concentration, turbidity measurements display a clear peptide threshold (saturation) concentration (Fig. 1d), which is characteristic of phase separation. It is worth noting that the phase boundary is almost superimposed with the electroneutrality line. In other words, phase separation occurs in the cationic regime, when the positive charges supplied by peptides outnumber the negative charges carried by RNA.
Figure 1: (a) Schematic representation of RRASLRRASL peptide (red) and polyU RNA (blue) forming condensates upon coulombic interaction. (b) A solution of condensates in a quartz cuvette exhibits a high level of turbidity. (c) Confocal microscopy image of condensates obtained with 1,000 µM of peptide including 0.5% of TAMRA-peptide, and 0.5 g.L-1 of RNA. (d) Contour map of turbidity measured at 500 nm as a function of the peptide and RNA concentrations. The red dashed line indicates electroneutrality.
The SAXS measurement of peptides reveals no sign of aggregation (Fig. LABEL:fig:SI-saxs_peptide) since the scattering curve at low $q$-values – $q$ being the wavenumber – is flat. Furthermore, the measured radius of gyration (0.9 nm) is consistent with the small number of amino acids (10). Similarly, the scattering curve of phosphorylated peptides reveals a radius of gyration of 0.84 nm (Fig. LABEL:fig:SI-saxs_peptide). PolyU RNA has $250\pm 80$ (s.d.) nucleotides according to mass photometry (Fig. LABEL:fig:SI-mass_photometry_polyU) and the radius of gyration is estimated to be 11.6 nm (Fig. LABEL:fig:SI-saxs_polyU). The Kratky representation of the RNA scattering curve displays an increase of the normalized intensity $(qR_{\mathrm{g}})^{2}(I/I_{0})$ – where $R_{\mathrm{g}}$ is the radius of gyration, $I$ the scattering intensity, and $I_{0}$ the forward scattering intensity – increases with $qR_{\mathrm{g}}$. This indicates a stretched conformation compared with a gaussian chain that will exhibit a flat Kratky representation – or alternatively, $I\sim q^{-2}$ –, as expected for a fully charged polyelectrolyte.
@@ 2.2 Passively-formed condensates grow by nucleated Brownian motion-induced coalescence
Due to the coulombic interactions underlying coacervation, growth is expected to start as early as a few milliseconds following the mixing of components, spanning length scales from several tens of nanometers to a few micrometers. Accordingly, TR-USAXS experiments on synchrotron source are carried out with a very large sample-to-detector distance of 31 m, and the rapid mixing is performed with a stopped-flow device that enables a dead time down to 2.5 ms. Figure 2a depicts scattering curves measured at different times for a peptide concentration of 1,000 µM (see Fig. LABEL:fig:SI-tr-saxs_passive for other peptide concentrations). The coacervation begins immediately after mixing and the remarkably flat Guinier region at low $q$-values is characteristic of the absence of interaction between condensates. Note that these measurements are repeatable from the first milliseconds up to nearly 10 min (Fig. LABEL:fig:SI-comp_growth_kinetics_1000uM).
@@
$I(t,q)\sim\displaystyle\int_{0}^{+\infty}\left[3\frac{\sin(qR)-qR\cos(qR)}{(qR)^{3}}\right]^{2}p(R;\langle R(t)\rangle,\Delta R(t))R^{6}dR$ (1)
@@
where $R$ is the condensate radius and $p(R;\langle R\rangle,\Delta R)$ the size distribution assumed here to be lognormal with mean radius $\langle R\rangle$ and width $\Delta R$, both varying with time. Figure 2b shows the time evolution of $\langle R\rangle$ inferred from the scattering curves at various peptide concentrations. Strikingly, the mean radius obeys a scaling law of the form $\langle R\rangle=(Kt)^{1/3}$ – where $K$ is a dynamic prefactor –, as early as the first few tens of milliseconds and this behavior persists over nearly four decades in time for the highest peptide concentrations. An exponent of $1/3$ during liquid-liquid phase separation is characteristic of either17, 19: (i) diffusion-limited Ostwald ripening where condensate-forming molecules migrate from small condensates to large condensates; or (ii) Brownian motion-induced coalescence during which condensates fuse to form increasingly larger condensates. In the latter case, the dynamic prefactor is expressed as $K\simeq 6k_{\mathrm{B}}T\phi/5\pi\eta$, with $k_{\mathrm{B}}$ the Botzmann constant, $T$ the temperature, $\phi$ the condensate volume fraction, and $\eta$ the solvent viscosity19. Whereas $K$ depends on the condensate volume fraction for the coalescence mechanism, it is only related to the properties of condensate-forming molecules such as the molar volume and the saturation concentration in the case of Ostwald ripening. The ratio of the measured dynamic prefactors for experiments performed with peptide concentrations of 1,000 and 2,000 µM while maintaining the peptide-to-RNA charge ratio constant, is estimated to be 1.6. Although a factor 2 is theoretically expected, the fact that $K$ increases with volume fraction suggests that coalescence mostly drives the growth kinetics.
@@
A quantity proportional to the number density of condensates $N$ can be deduced from the forward scattering intensity $I_{0}\equiv I(q\rightarrow 0)$ and the size distribution at each time since $I_{0}\sim N\langle V^{2}\rangle\sim N\langle R^{6}\rangle$ with $V$ the condensate volume35. Figure 2c depicts this quantity versus time and demonstrates that $N$ scales as $t^{-1}$, collapsing onto a single curve at late times irrespective of the peptide concentration. This behavior is consistent with a coalescence mechanism: the volume fraction $\phi$ remains constant over time, and applying the scaling law for coalescence, $R^{3}\sim\phi t\sim NR^{3}t$, leads to $N\sim t^{-1}$, with a prefactor that is independent of $\phi$.
@@
Additionally, Figure 2b illustrates the nucleation preceding the condensate growth. As the peptide concentration approaches the saturation concentration ($\approx$400 µM), an increasingly long induction time is observed before a sufficient signal-to-noise ratio on the scattering intensity permits to infer condensate size (see the 400 and 600 µM curves on Fig. 2b). A close examination of the SAXS curves at 400 µM reveals no detectable evolution until 32 s, while within the following seconds, the scattering intensity dramatically increases. We hypothesize that prior to nucleation, each isolated RNA chain is decorated with bound peptides. This interpretation is supported by a comparison between the average scattering intensity measured prior to nucleation and the sum of the individual contributions from peptides and RNA. As the former exceeds the latter (Fig. LABEL:fig:SI-supersaturation), peptides are indeed bound to RNA. From the forward scattering intensity, we estimate that approximately 30 peptides are associated with each RNA chain (see Supporting Information). Furthermore, fitting the average scattering intensity using the Fisher–Burford model (Fig. LABEL:fig:SI-supersaturation) 36, which describes simple fractal aggregates, yields a mass fractal dimension of 1.4. This value indicates that the clusters adopt an extended, chain-like conformation, with an overall size ($R_{\mathrm{g}}=13\ \text{nm}$) comparable to that of bare RNA ($R_{\mathrm{g}}=11.6\ \text{nm}$).
@@
To get deeper insights into the early events at the molecular level, we perform coarse-grained molecular dynamics (MD) simulations with implicit solvent. The numerical model is able to reproduce a saturation concentration of 350 µM (Fig. LABEL:fig:SI-MD_phase_diagram), that is, within the range of values found experimentally (Fig. 1d; $\approx$400 µM). Peptides and RNA are initially mixed in a simulation box, and Figure 3 depicts the time evolution of the number of peptides contained in the largest peptide-RNA cluster. Very rapidly, all the peptides bind on RNA, which are thus decorated with an average of 50 peptides. Interestingly, the total number of bound peptides barely varies afterwards (Fig.LABEL:fig:SI-MD_largest_cluster_allc), confirming the role of peptide-decorated RNA as subunits. The peptide-RNA clusters further grow in a stepwise manner by coalescence. This is illustrated on Fig.3 wherein the steps observed in the number of peptides in the largest cluster become higher over the course of the simulation. In particular, snapshots 4 and 5 show two clusters comprising 400 and 200 peptides, respectively, that is, 8 and 4 RNA chains (Fig. LABEL:fig:SI-MD_largest_cluster_allb), respectively, fusing into a single cluster. Note that fusion events are increasingly rare with the simulation time steps as the diffusion of larger clusters is slower.
@@
At the macroscopic level, mean radii exceeding a few micrometers cannot be probed reliably even by ultra-small-angle X-ray scattering. Therefore, we carry out confocal microscopy imaging by using again 0.5% of fluorescent TAMRA-peptide. Figure 4a depicts a low magnification view of condensates at a peptide concentration of 1,000 µM, 25 min after mixing with RNA. The radius distribution, shown in Fig. 4b, resembles a lognormal distribution, as expected from a coalescence mechanism 19. No macroscopic phase separation is observed, most likely as a result of coulombic repulsion that self-regulates the growth mechanism of the condensates 21.

### v2

@@ Abstract
Living cells exhibit a complex organization comprising numerous compartments, among which are RNA- and protein-rich membraneless, liquid-like organelles known as biomolecular condensates. Energy-consuming processes regulate their formation and dissolution, with (de-)phosphorylation by specific enzymes being among the most commonly involved reactions. By employing a model system consisting of a phosphorylatable peptide and homopolymeric RNA, we elucidate how enzymatic activity modulates the growth kinetics and alters the local structure of biomolecular condensates. Under passive condition, time-resolved ultra-small-angle X-ray scattering with synchrotron source reveals a nucleation-driven coalescence mechanism maintained over four decades in time, similar to the coarsening of simple binary fluid mixtures. Coarse-grained molecular dynamics simulations show that peptide-decorated RNA chains assembled shortly after mixing constitute the relevant subunits. In contrast, actively-formed condensates initially display a local mass fractal structure, which gradually matures upon enzymatic activity before condensates undergo coalescence. Both types of condensate eventually reach a steady state but fluorescence recovery after photobleaching indicates a peptide diffusivity twice higher in actively-formed condensates consistent with their loosely-packed local structure. We expect multiscale, integrative approaches implemented with model systems to link effectively the functional properties of membraneless organelles to their formation and dissolution kinetics as regulated by cellular active processes.
@@ 1 Introduction
Membraneless organelles are intracellular compartments devoid of lipid membrane, consisting of multiple proteins and possibly nucleic acids (DNA and/or RNA), and formed by liquid-liquid phase separation 1, 2, 3, 4. These biomolecular condensates perform essential biological functions by sequestering specific factors and promoting biochemical reactions following the needs of the cell 5. They also serve as protocells for the study of the emergence of functional cellular life or as catalytic microcompartments 6, 7. Many condensates found in eucaryotic cells contain RNA and are involved in gene expression, such as nucleoli 8, P granules 9, stress granules 10 or Cajal bodies 11 to name a few. Protein-RNA condensates may also play a role in pathological processes 12, for example by facilitating the transcription of viral RNA during SARS-Cov-2 replication 13. Membraneless organelles are relatively short-lived, in that they are formed to perform specific functions and are dissolved after the completion of these tasks. Such a flexibility is maintained through energy-consuming enzymatic processes. In particular, various post-translational modifications have been shown to regulate intracellular phase separation, including phosphorylation 14, methylation 15, acetylation and combinations thereof 16.
@@
Phase separation is a phenomenon by which a homogeneous material, spontaneously or not, separates into two or more distinct domains (phases) with different physical and/or chemical properties. Extensive experimental and theoretical efforts have been devoted to elucidating the phase-separation kinetics of binary systems 17, including fluid mixtures relevant to biomolecular condensates, under both passive (spontaneous) and active (chemically driven) conditions 18, 19, 20. A major limitation in these investigations lies in the lack of information across length and time scales spanning several orders of magnitude. Indeed, biomolecular condensates are mostly micrometer-sized and therefore, fluorescence microscopy has become the standard technique to monitor their life cycle, especially within cells 8. Quite recently, holographic microscopy has demonstrated its ability to monitor the growth of PopZ-cation condensates without the need for protein labeling 21. In all these cases, submicrometer-sized condensates formed along the growth pathway are overlooked, and events occurring on subsecond timescales often remain inaccessible. Conversely, coarse-grained molecular dynamics simulations can provide a wealth of information at the molecular level 22, and a few phase-separated protein-RNA condensates have been elucidated 23, 24. Nonetheless, coarse-grained models face substantial challenges in quantitatively reproducing the thermodynamic properties of biomolecular condensates. In addition, although they allow the simulation of relatively large supramolecular systems, the accessible timescales remain limited to a few microseconds at most.
@@
To bridge these multiple scales, time-resolved small-angle X-ray scattering (TR-SAXS) using synchrotron source has emerged as a state-of-the-art technique, enabling to track self-assembly processes with millisecond temporal resolution and spatial sensitivity ranging from a few to several hundred nanometers 25, 26, 27, 28, 29. Within the context of liquid-liquid phase separation, the assembly kinetics of a prion-like domain has revealed the existence of a double nucleation barrier to overcome before the condensate growth can proceed 30. With the advent of time-resolved ultra-small-angle X-ray scattering (TR-USAXS) 31 available at a very few synchrotron facilities worldwide, the capabilities of X-ray scattering have extended to the micrometer scale. For example, the spinodal decomposition of bovine serum albumin induced by YCl3 was probed during the first two minutes following the mixing of components 32. Likewise, coacervates composed of oppositely-charged polyelectrolytes were found to change from mass fractal to surface fractal morphologies during their growth driven by coalescence 33. Nevertheless, how enzymatic activity regulates the growth mechanism, the local structure, and the morphology of biomolecular condensates – compared to conditions lacking active processes – remains experimentally unaddressed.
@@
To fill this knowledge gap, we employ a model system of biomolecular condensates whose formation is triggered by the dephosphorylation of serine residues, thereby mimicking the enzymatic regulation encountered in intracellular membraneless organelles 34. This system consists of an arginine-rich peptide (RRASLRRASL) containing phosphorylatable serines, associated with poly(uridylic acid) (polyU) RNA. Condensates form spontaneously when the serines bear positive charges – a condition hereafter referred to as passive – or are enzymatically induced by lambda phosphatase (LPP), which removes phosphate groups from initially phosphorylated serines, defining the active condition. This flexible system enables the comparison between condensate formation mechanisms under passive and active conditions. We therefore combine TR-USAXS experiments performed at varying peptide and enzyme concentrations, with confocal microscopy and coarse-grained molecular dynamics simulations to capture these mechanisms across a wide range of length- and timescales.
@@ 2.1 Phase separation occurs in the cationic regime
Upon mixing RRASLRRASL peptide with anionic polyU RNA (Fig. 1a) in 4 mM MgCl2, 2.25 mM ATP, 1.6 mM dithiothreitol, 50 mM HEPES pH 7.4 at 37 °C, biomolecular condensates in the form of liquid-like droplets are spontaneously made by coacervation 34. A direct evidence for the presence of micrometer-sized droplets is given by the turbid aspect of the solution (Fig. 1b). Confocal microscopy images performed with 0.5% of fluorescent TAMRA-peptide confirm the spherical shape and the micrometer-scale size of the condensates (Fig. 1c). At a given RNA concentration, turbidity measurements display a clear peptide threshold (saturation) concentration (Fig. 1d), which is characteristic of phase separation. It is worth noting that the phase boundary is almost superimposed with the electroneutrality line. Phase separation occurs in the cationic regime, when the total positive charges supplied by the peptides outnumber the total negative charges carried by RNA.
Figure 1: (a) Schematic representation of RRASLRRASL peptide (red) and polyU RNA (blue) forming condensates upon Coulomb interaction. (b) A solution of condensates in a quartz cuvette exhibits a high level of turbidity. (c) Confocal microscopy image of condensates obtained with 1,000 µM of peptide including 0.5% of TAMRA-peptide, and 0.5 g.L-1 of RNA. (d) Contour map of turbidity measured at 500 nm as a function of peptide and RNA concentrations. The red dashed line indicates electroneutrality.
The SAXS measurement of peptides reveals no sign of aggregation (Fig. LABEL:fig:SI-saxs_peptide) given that the scattering curve at low $q$-values – $q$ being the magnitude of the scattering wavevector – is flat, which means that the Guinier plateau is reached. Furthermore, the measured radius of gyration ($R_{\mathrm{g}}=\text{0.9\ nm}$) is consistent with the small number of amino acids (ten). Likewise, the scattering curve of phosphorylated peptides reveals a radius of gyration of 0.84 nm (Fig. LABEL:fig:SI-saxs_peptide). PolyU RNA has $250\pm 80$ (s.d.) nucleotides according to mass photometry (Fig. LABEL:fig:SI-mass_photometry_polyU) and its radius of gyration is estimated to be 11.6 nm (Fig. LABEL:fig:SI-saxs_polyU). The Kratky representation of the RNA scattering curve displays an increase of the normalized intensity $(qR_{\mathrm{g}})^{2}(I/I_{0})$ at high $qR_{\mathrm{g}}$ – where $I$ stands for the scattering intensity and $I_{0}$ is the forward scattering intensity –, which indicates a stretched conformation as expected for a fully charged polyelectrolyte.
@@ 2.2 Passively-formed condensates grow by nucleated Brownian motion-induced coalescence
Due to the Coulomb interactions underlying coacervation, growth is expected to start as early as a few milliseconds following the mixing of components, spanning lengthscales from several tens of nanometers to a few micrometers. Accordingly, TR-USAXS experiments on synchrotron source are carried out with a very large sample-to-detector distance of 31 m, and the rapid mixing is performed with a stopped-flow device that enables a dead time as short as 2.5 ms. Figure 2a depicts scattering curves measured at different times for a peptide concentration of 1,000 µM (see Fig. LABEL:fig:SI-tr-saxs_passive for other peptide concentrations). The coacervation begins immediately after mixing and the remarkably flat Guinier plateau at low $q$-values is characteristic of the absence of interaction between condensates. Note that these measurements are repeatable from the first milliseconds up to nearly 10 min (Fig. LABEL:fig:SI-comp_growth_kinetics_1000uM).
@@
$I(t,q)=\frac{I_{0}(t)}{\langle R^{6}(t)\rangle}\displaystyle\int_{0}^{+\infty}\left[3\frac{\sin(qR)-qR\cos(qR)}{(qR)^{3}}\right]^{2}p(R;\langle R(t)\rangle,\Delta R(t))R^{6}dR+C_{\mathrm{bkg}}$ (1)
@@
where $R$ is the condensate radius and $p(R;\langle R(t)\rangle,\Delta R(t))$ the size distribution assumed here to be lognormal with mean radius $\langle R(t)\rangle$ and width $\Delta R(t)$. $C_{\mathrm{bkg}}$ is a constant accounting for the background. The three fitting parameters at each time step are therefore $I_{0}$, $\langle R\rangle$ and $\Delta R$, $\langle R^{6}\rangle$ being computed with the size distribution. Figure 2b shows the time evolution of $\langle R\rangle$ inferred from the scattering curves at various peptide concentrations. Strikingly, the mean radius obeys a scaling law of the form $\langle R\rangle=(Kt)^{1/3}$ – where $K$ is a dynamic prefactor –, as early as the first few tens of milliseconds, and this behavior persists over nearly four decades in time for the highest peptide concentrations. An exponent of $1/3$ during liquid-liquid phase separation is the hallmark of either17, 19: (i) diffusion-limited Ostwald ripening where condensate-forming molecules migrate from small condensates to large condensates; or (ii) Brownian motion-induced coalescence during which condensates fuse to form increasingly larger condensates. In the latter case, the dynamic prefactor is expressed as $K\simeq 6k_{\mathrm{B}}T\phi/5\pi\eta$, with $k_{\mathrm{B}}$ the Botzmann constant, $T$ the temperature, $\phi$ the condensate volume fraction, and $\eta$ the solvent viscosity19. Whereas $K$ depends on the condensate volume fraction for the coalescence mechanism, it is only related to the properties of condensate-forming molecules such as the molar volume and the saturation concentration in the case of Ostwald ripening. The ratio of the measured dynamic prefactors for experiments performed with peptide concentrations of 1,000 and 2,000 µM while maintaining the peptide-to-RNA charge ratio constant, is estimated to be 1.6. Although a factor 2 is theoretically expected, the fact that $K$ increases with volume fraction suggests that coalescence mostly drives the growth kinetics.
@@
A quantity proportional to the number density of condensates $N$ can be deduced from the forward scattering intensity $I_{0}\equiv I(q\rightarrow 0)$ and the size distribution at each time since $I_{0}\sim N\langle V^{2}\rangle\sim N\langle R^{6}\rangle$ with $V$ the condensate volume 35. Figure 2c depicts this quantity versus time and demonstrates that $N$ scales as $t^{-1}$, collapsing onto a single curve at late times irrespective of the peptide concentration. This behavior is consistent with a coalescence mechanism: the volume fraction $\phi$ remains constant over time, and applying the scaling law for coalescence, $R^{3}\sim\phi t\sim NR^{3}t$, leads to $N\sim t^{-1}$, with a prefactor that is indeed independent of $\phi$.
@@
Additionally, Figure 2b illustrates the nucleation preceding the condensate growth. As the peptide concentration approaches the saturation concentration ($\approx$400 µM), an increasingly long induction time is observed before a sufficient signal-to-noise ratio on the scattering intensity enables to infer condensate size (see the 400 and 600 µM curves on Fig. 2b). A close examination of the SAXS curves at 400 µM reveals no detectable evolution until $t=32\text{ s}$, while within the following seconds, the scattering intensity dramatically increases. We hypothesize that prior to nucleation, each isolated RNA chain is decorated with bound peptides. This interpretation is supported by a comparison between the average scattering intensity measured prior to nucleation and the sum of the individual contributions from peptides and RNA. As the former exceeds the latter (Fig. LABEL:fig:SI-supersaturation), peptides are indeed bound to RNA. From the forward scattering intensity, we estimate that approximately 30 peptides are associated with each RNA chain (see Supporting Information). Furthermore, fitting the average scattering intensity using a Fisher–Burford model (Fig. LABEL:fig:SI-supersaturation) 36, which describes simple fractal aggregates, yields a mass fractal dimension of 1.4. This value indicates that the clusters adopt an extended, chain-like conformation, with a size ($R_{\mathrm{g}}=13\ \text{nm}$) comparable to that of bare RNA ($R_{\mathrm{g}}=11.6\ \text{nm}$).
@@
To get deeper insights into the early events at the molecular level, we perform coarse-grained molecular dynamics (MD) simulations with implicit solvent by using a recently released CALVADOS RNA force field 37. The numerical model is able to reproduce a saturation concentration of 350 µM (Fig. LABEL:fig:SI-computed_phase_diagram), that is, within the range of values found experimentally (Fig. 1d; $\simeq$400 µM). Peptides and RNA are initially mixed in a simulation box, and Figure 3 depicts the time evolution of the number of peptides contained in the largest peptide-RNA cluster. Very rapidly, all the peptides bind on RNA, which are thus decorated with an average of 50 peptides. Remarkably, the total number of bound peptides barely varies afterwards (Fig.LABEL:fig:SI-MD_largest_cluster_allc), confirming the role of peptide-decorated RNA as stable subunits. The peptide-RNA clusters further grow in a stepwise manner by coalescence. This is illustrated on Fig.3 wherein the steps observed in the number of peptides in the largest cluster become higher over the course of the simulation. In particular, snapshots 4 and 5 show two clusters comprising 400 and 200 peptides, respectively, that is, 8 and 4 RNA chains (Fig. LABEL:fig:SI-MD_largest_cluster_allb), respectively, merging into a single cluster. Note that fusion events are increasingly rare with the simulation time steps as the diffusion of larger clusters is much slower.
@@
At a macroscopic level, condensates exceeding several micrometers in size cannot be probed even by USAXS. Therefore, we carry out confocal microscopy imaging by using again 0.5% of fluorescent TAMRA-peptide. Figure 4a depicts a low magnification view of condensates at a peptide concentration of 1,000 µM, 25 min after mixing with RNA. The radius distribution, shown in Fig. 4b, resembles a lognormal distribution, as expected from a coalescence mechanism 19. No macroscopic phase separation is observed, most likely as a result of Coulomb repulsion that self-regulates the growth of condensates 21.